If Drink A contains 81% caffeine and Drink B contains 69% caffeine. And you want a mixture of both solutions to to create 150 gallons of solution which contains 71% caffeine.
Which system of equations should you use to solve this problem:
x+y=150
81x+69y=106.5
x+y=150
0.81x+0.69y=106.5
x+y=106.5
81x+69y=150
x+y=150
81x+69y=71

Answer by ikleyn(52786)

.
If Drink A contains 81% caffeine and Drink B contains 69% caffeine. And you want a mixture of both solutions
to create 150 gallons of solution which contains 71% caffeine.

Which system of equations should you use to solve this problem:

x+y=150
81x+69y=106.5

x+y=150
0.81x+0.69y=106.5

x+y=106.5
81x+69y=150

x+y=150
81x+69y=71
~~~~~~~~~~~~~~~~~~


                My     ANSWER

    x +     y = 150
0.81x + 0.69y = 106.5.

Second system of equations from your list.

--------------

If you want   to  LEARN   on how to setup the governing equations for the mixture problems,
read these introductory lessons
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Mixture problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.